OpenStudy (anonymous):

What is the third root of -216?

OpenStudy (johnweldon1993):

What number...times itself 3 times = -216? hint * -5^3 = -5 * -5 * -5 = -125 -7^3 = -7 * -7 * -7 = -343 so you answer is somewhere between there...can you figure out what it is?

OpenStudy (anonymous):

do you know how to write the third root as an exponent, for example, the sqrt(4) = 4^(1/2)

OpenStudy (anonymous):

7 @johnweldon1993

OpenStudy (whpalmer4):

7*7 = 49 49*7 = something other than 216

OpenStudy (anonymous):

I AM SO NOT GOOD AT THESE @whpalmer4

OpenStudy (whpalmer4):

well, 7 is too big: 7*7*7 = 49*7 = 280+63=343 how about 5? 5*5*5 = 25*5 = 125, too small how about 6? 6*6*6 = 36*6 = 180+36=216 Eureka!

OpenStudy (whpalmer4):

and we need it to equal -216, so -6*-6*-6 = 36*-6 = -216